I am asked to create a quadratic function, y=f(x), that satisfies the following conditions:
f(7)=2, f(-5)=-10, and y-intercept is (0,6)
NO clue where to begin. I know I could do something if I had the x-intercepts. But what I start with is:
6=ax^2+bx+c , but then I come up with 6=c and I feel like I just need a starting point.
I was not tought this in my current class, supposedly this is suppose to be prerequisite homework...
Thanks for any help...Kate
Answer by Alan3354(69443) 
 f(x) = ax^2 + bx + c
Find a, b & c is the problem.
--------------
c = 6, the y-intercept, because
f(0) = 6
------------
f(7) = 49a + 7b + 6 = 2
f(-5) = 25a - 5b + 6 = -10
Now it's 2 eqns with 2 variables, a & b.
Can you finish it?
-----------------
I get f(x) = -11x^2/35 + 57x/35 + 6
| |
|